Problem 6 Medical Corporation of America (MCA) has a current stock price of $36, and its last dividend (D) was $2.40. In view of MCAs strong financial position, its required rate of return is 12 percent. If MCAs dividends are expected to grow at a constant rate in the future, what is the firms expected stock price in five years?
Answer #1

Stock price = $36

Required rate of return = 12%

Dividend just paid = $2.40

From Gordon growth model,

Price = Dividend next year / (Required return – growth)

36 = 2.40 * (1+g)/ (0.12 – g)

Solving for g, we get growth rate = 5%

Considering growth rate in earnings to be 5 % per year, Dividend payment in five years will be

Dividend just paid * (1.05)^5 = 2.40 * 1.27 = $3.06

Stock price in five years = Dividend in 6th year/ Required return – Growth)

                                            = 3.06 * 1.05 / (0.12-0.05) = 3.21/0.07 = $45.85